OpenStudy (anonymous):

Find a polynomial function f(n) such that f(1), f(2), ... , f(8) is the following sequence. 4, 10, 16, 22, 28, 34, 40, 46

OpenStudy (anonymous):

Its Discrete Math

OpenStudy (anonymous):

I got 6n-4 but was not correct

OpenStudy (raden):

looks it is a linear function. let f(n) = an + b when f(1) = 4, it means a(1)+b = 4 or a + b = 4 .... (1st equation) and f(2) = 10, it means a(2) + b = 10, or 2a + b = 10 ..... (2nd equation) by elimination methode, subtract both equations get a + b = 4 2a + b = 10 ------------ (-) -a = -6 a = 6 subtitute a = 6 into one of the equation above, let take 1st. a+b = 4 6 + b = 4 b = 4 - 6 b = -2 see that f(n) = an + b. distribute the value of a and b to f(n), get f(n) = 6n-2

OpenStudy (anonymous):

Are you sure its 6n-2 tho?

OpenStudy (raden):

we can cross-check, if f(n) = 6n -2 as the correct answer or not f(n) = 6n - 2 check the values given above : f(1) = 6(1) - 2 = 4 (true) f(2) = 6(2) - 2 = 10 (true) f(3) = 6(3) - 2 = 16 (true) f(4) = 6(4) - 2 = 22 (true) f(5) = 6(5) - 2 = 28 (true) f(6) = 6(6) - 2 = 34 (true) f(7) = 6(7) - 2 = 40 (true) f(8) = 6(8) - 2 = 46 (true) so, dont worry. that's right!

OpenStudy (anonymous):

Okay okay thank you!

OpenStudy (raden):

you're welcome ^_^